A buck converter is supplied from a 50V battery source. Given L=400μH, C=100μF, R=20Ω, f=20KHz and D=0.4 Calculate:
(a) output voltage
(b) maximum and minimum inductor current,
(c) output voltage ripple,
(d) What are the effects on the voltage ripple if the switching frequency is doubled?
Answer #1

400 UH 00012N + I DEDOL f= 206 1 50V / A VE TE sw C E20v I max = I +A IL 80 - 100uF / 1200 Uoutput voltage vipple cusxewt AIL